Assuming the population has a normal distribution. A sample of 25 randomly selected students has a mean test score of 81.5 woth a stardard deviation of 10.2.
A. costruct a 95% confidence interval for the population mean.(show all five steps please)
BInterpret the confidence in the context of this problem.
CInterpret the confidence in the context of this problem..
IT WOULD BE GREAT IT YOU COULD HELP ME THANK YOU SO MUCH FOR YOUR TIME!!!!

Answer by stanbon(75887)

Assuming the population has a normal distribution. A sample of 25 randomly selected students has a mean test score of 81.5 with a standard deviation of 10.2.
A. contruct a 95% confidence interval for the population mean.(show all five steps please)
sample mean: 81.5
standard error: E = t*s/sqrt(25) = 2.064*10.2/sqrt(25) = 4.21
95% CI: 81.5-4.21 < u < 81.5+4.21
95% CI: 77.29 < u < 85.71
--------------------------
B. Interpret the confidence in the context of this problem.
We are 95% confident the population mean is between 77.29 and 85.71
==================
Cheers,
Stan H.